# State-Constrained Offline Reinforcement Learning

### v1

@@ 4.2 An illustrative example: maze navigation
To elucidate the advantages of SCQL in comparison to BCQL, we explore a simple maze environment backed by a dataset with few trajectories. The maze is visualised as a $10$ by $10$ grid. The available states are shown as white squares, while the red squares represent the maze walls. Each state in this environment is denoted by the $(x,y)$ coordinates within the maze, and the agent’s available actions comprise of simple movements: left, right, up, and down. With each movement, the agent incurs a reward penalty of a small negative value, but upon successfully navigating to the gold star, it is rewarded with a large positive value. Consequently, the agent’s primary objective becomes devising the most direct route to the star. In this deterministic environment, the outcomes of transitioning to a state and executing an action are congruent, meaning Eq. (1) and (2) mirror each other. This allows for a direct comparison between BCQL and SCQL in terms of their performance and ability to leverage limited datasets.
@@
Figure 1: Comparison of BCQL and SCQL methods on a simple maze environment. (a) The maze is a 10 by 10 grid where the co-ordinate values $(x,y)$ represent the state. The high reward region is represented by a star and maze walls are represented by red grid squares. The dataset is made of 4 trajectories represented by the dotted lines where the open circle is the starting state and the closed circle is the final state. (b) The final policy when applying BCQL to the dataset. (c) The final policy when applying SCQL to the dataset.
Figure 1a provides a visual representation of the maze and accompanying dataset. Here, trajectories start at the white circle and end at the black one. Despite its simplicity, the dataset contains a scant number of states wherein multiple actions have been taken. Within this specific framework, the notion of state reachability is straightforward: any state that is just one block away from the current state is considered reachable. Formally the state reachability is, $\mathcal{SR}_{\mathcal{M}}((x,y))=\{(x+1,y),(x-1,y),(x,y+1),(x,y-1)\}$. In other words, a state is considered reachable from the current state if it is one block away in any of the four cardinal directions (left, right, up, or down). This simple definition of state reachability is sufficient for demonstrating the advantages of SCQL in this illustrative example.
Applying BCQL with training as per Eq. (3) and policy extraction as per
$\pi(s)=\max_{a\text{ s.t. }(s,a)\in\mathcal{D}}Q(s,a),$
we obtain the policy depicted in Figure 1b, where arrows indicate policy direction. Conversely, for SCQL, training via Eq. (6) and policy training through Eq. (7) yields the policy in Figure 1c. For clarity, if no optimal movement can be found in the state, due to no information in the dataset, it is left blank. SCQL is able to leverage the sparse dataset more efficiently, enabling the agent to reach the gold star from any dataset state position. In contrast, BCQL succeeds in only 11 of the 57 unique dataset states, this is equivalent to behavioural cloning’s performance. SCQL’s advantage stems from its ability to leverage state adjacency information to stitch together optimal trajectories, whereas BCQL relies solely on previously observed state-action pairs, limiting its performance in datasets with a small number of trajectories.
@@ 5 Practical implementation: StaCQ
In this section, we introduce StaCQ (State-Constrained deep QS-learning), our particular deep learning implementation of SCQL. We present a policy constraint method that regularises the policy towards the best reachable next state in the dataset. Our specific implementation learns a QS-function that is aligned with the underlying theory. However, state-constrained methods could be adapted from the theory in many ways, analogous to many techniques that modify the batch-constraint for enhancing BCQ [22]. For example, alternative state-constrained methods could incorporate different policy regularisation techniques, such as those based on divergence measures or constraint violation penalties, or they could explore different ways of estimating state reachability based on the available data and domain knowledge.
@@ 5.1 Estimating state reachability
Central to the state-constrained approach is the concept of state reachability, as per Definition 4.1. Because the environments used in our benchmarks do not provide this information, we need to estimate reachability from the dataset. In our implementation, we consider $\hat{s}^{\prime}$ reachable to $s$ if we can predict the action that reaches $\hat{s}^{\prime}$. This entails the use of two key models: a forward dynamics model, $f_{\omega_{1}}(s,a)$, and an inverse dynamics model, $I_{\omega_{2}}(s,s^{\prime})$. Both of these models are framed as neural networks and are trained via supervised learning using triplets of form $(s,a,s^{\prime})\sim\mathcal{D}$. The loss function for the forward model is given by
@@
Consistent with prior research [9, 13, 28, 4, 59, 58], we adopt an ensemble strategy. Each model within the ensemble possesses a distinct set of parameters. The ensemble’s final predictions for $s^{\prime}$ and $a$ are computed by taking the average of their respective outputs. It should be noted that these are very simple models and adding more complexity such as [62] will greatly improve model prediction and thus state reachability prediction.
Using these models, we propose an estimate of state reachability, denoted by $\widehat{\mathcal{SR}}_{\mathcal{M}}$, as:
@@
This criterion suggests that $\hat{s}^{\prime}$ is reachable from $s$ only if an action can be predicted that transitions the agent from $s$ to $\hat{s}^{\prime}$. It is important to note that $\epsilon$ is a small positive value to account for possible model inaccuracies. In Appendix B we detail a way of reducing the complexity of this calculation.
@@ 5.2 StaCQ
StaCQ makes use of an actor-critic framework; the critic, i.e. the $Q(s,s^{\prime})$ value of transitioning from $s$ to $s^{\prime}$, is trained by minimising the mean square error (MSE) between predicted and true QS-values. Following the theory, the QS-values are updated on all pairs of states and reachable next states:
$\mathcal{L}_{\theta}=\mathbb{E}_{\begin{subarray}{c}s\sim\mathcal{D}\\ \hat{s}^{\prime}\in\widehat{\mathcal{SR}}_{\mathcal{M}}(s)\end{subarray}}\Biggl{[}\Biggl{(}r(s,\hat{s}^{\prime})+\gamma Q_{\theta^{\prime}}(\hat{s}^{\prime},f_{\omega_{1}}(\hat{s}^{\prime},\pi_{\phi}(\hat{s}^{\prime})))-Q_{\theta}(s,\hat{s}^{\prime})\Biggr{)}^{2}\Biggr{]}.$ (11)
In this equation, the actor produces an action, $\pi_{\phi}(s^{\prime})$, which is input into the forward model $f_{\omega_{1}}$, before being evaluated in the QS-function. The specific reward for the reachable pair $(s,\hat{s}^{\prime})$ is potentially unseen in the dataset, therefore if the reward function is unknown it must be approximated using a neural network. In a manner akin to the forward and inverse dynamics models, this can be achieved using supervised learning by reducing the mean square error (MSE) between the predicted and actual rewards:
@@
Past research indicates that this type of reward model is proficient in estimating rewards for previously unseen transitions [26]. In Eq. (11), $\theta^{\prime}$ represents the parameters for a target Q-value which are incrementally updated towards $\theta$: $\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$, with $\tau$ being the soft update coefficient.
To train the actor, we seek to maximise the current QS-values while staying close to the best next state in the dataset, therefore our loss to minimise is:
$\mathcal{L}_{\phi}=\mathbb{E}_{s\sim\mathcal{D}}\Bigl{[}\lambda Q_{\theta}(s,f_{\omega_{1}}(s,\pi_{\phi}(s)))+(f_{\omega_{1}}(s,\pi_{\phi}(s)))-\hat{s}^{\prime})^{2}\Bigr{]}.$ (13)
In this equation we have a $\lambda$ hyperparameter which provides our level of regularisation. The $\hat{s}^{\prime}$ is the best reachable next state, from $s$, according to the QS-value,
$\hat{s}^{\prime}=\arg\max_{s^{\prime}\in\widehat{\mathcal{SR}}_{\mathcal{M}}(s)}Q_{\theta}(s,s^{\prime}).$ (14)
@@
3:  Pre-train $f_{\omega_{1}}$ $I_{\omega_{2}}$: Eqs.(8) & (9)
@@
6:     Optimise reward function: Eq. (12)
7:     Optimise QS-value: Eq. (11)
8:     Optimise policy: Eq. (13)
9:     Update target networks: $\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$.
@@
This is therefore a similar policy extraction method to TD3+BC [21], where both methods use a BC regulariser. However, StaCQ is able to utilise the state reachability metric enabling the maximisation to be taken over more states, whereas TD3+BC only has a single state-action pair. Due to the forward model being fixed, a small amount of Gaussian noise is added to the policy action before being input into the model. This creates a more robust policy by ensuring the policy does not exploit the forward model. The full algorithm is shown in Algorithm 1. Notably, all three constituent models are trained using supervised learning, making them straightforward sub-processes within this architecture.
@@ 6 Experimental results
We evaluate StaCQ against several model-free and model-based baselines, on the D4RL benchmarking datasets from the OpenAI Mujoco tasks [52, 20]. The model-free baselines we compare to are BCQ [22], TD3+BC [21] and IQL [32]. BCQ is the most fundamental batch-constrained offline RL method which many future methods are based from and is thus the most direct comparison theory-wise. TD3+BC is the most similar method to StaCQ implementation wise as both techniques use a BC-style regularisation process. IQL is a current state-of-the-art (SOTA) model-free method which also provides a useful comparison. The model-based baselines we compare to are MBTS [26], Diffuser [27] and RAMBO [42]. MBTS conducts a data augmentation strategy that uses a different methodology of state reachability to StaCQ. Diffuser is a planning based method that uses a diffusion model, thus a more complex dynamics model than StaCQ. RAMBO is a SOTA model-based method that performs rollouts using a learned pessimistic dynamics model of the environment. Although StaCQ does not use models for planning nor rollouts, we include these model-based algorithms for comparison purposes. We have also devised a one-step [8] version of StaCQ, detailed in Appendix D, showing the flexibility of the general state-constrained framework.
@@
Hopper Rand $6.2$ $7.6$ $8.5$ $\mathbf{21.6}$ $9.1\pm 0.9$ $7.5\pm 0.4$
Med-Rep $22.5$ $51.0$ $60.9$ $\mathbf{94.7}$ $50.2$ $93.6$ $\mathbf{96.6}$ $\mathbf{99.1\pm 0.7}$ $\mathbf{99.4\pm 0.6}$
Med $56.8$ $60.9$ $59.3$ $66.3$ $64.3$ $74.3$ $92.8$ $\mathbf{101.8\pm 0.3}$ $93.3\pm 3.2$
Med-Exp $54.2$ $85.9$ $98.0$ $91.5$ $94.8$ $103.3$ $83.3$ $\mathbf{111.2\pm 1.7}$ $92.1\pm 9.2$
@@
Med $39.4$ $73.7$ $83.7$ $78.3$ $78.8$ $79.6$ $86.9$ $\mathbf{93.1\pm 1.2}$ $85.7\pm 10.4$
Med-Exp $90.5$ $94.5$ $\mathbf{110.1}$ $\mathbf{109.6}$ $108.8$ $106.9$ $68.3$ $\mathbf{114.8\pm 1.5}$ $107.6\pm 8.5$

### v2

@@ 3.2 An illustrative example: maze navigation
To elucidate the advantages of SCQL in comparison to BCQL, we explore a simple maze environment backed by a dataset with few trajectories. This pedagogical illustration aims to show how SCQL utilises the concept of state reachability - which enables improved performance over BCQL (which has no mechanism to utilise state reachability). Figure 2 shows the full maze environment, visualised as a $10$ by $10$ grid. The available states are shown as white squares, while the red squares represent the maze walls. Each state in this environment is denoted by the $(x,y)$ coordinates within the maze, and the agent’s available actions comprise of simple movements: left, right, up, and down. With each movement, the agent incurs a reward penalty of a small negative value ($r_{\text{pen}}=-0.1$), but upon successfully navigating to the gold star, it is rewarded with a large positive value ($r_{\text{goal}}=10$). Consequently, the agent’s primary objective becomes devising the most direct route to the star. In this deterministic environment, the outcomes of transitioning to a state and executing an action are congruent, meaning Eq. equation 1 and equation 2 mirror each other. This allows for a direct comparison between BCQL and SCQL in terms of their performance and ability to leverage limited datasets. This experiment converges after $100$ steps with a learning rate of $\alpha=0.25$ and a discount factor of $\gamma=0.99$.
@@
Figure 2: Comparison of BCQL and SCQL methods on a simple maze environment. (a) The maze is a 10 by 10 grid where the coordinate values $(x,y)$ represent the state. The high reward region is represented by a star and maze walls are represented by red grid squares. The dataset is made of 4 trajectories represented by the dotted lines where the white circle is the starting state and the black circle is the final state. (b) The final policy when applying BCQL to the dataset. (c) The final policy when applying SCQL to the dataset.
Figure 2a provides a visual representation of the maze and accompanying dataset. Here, trajectories start at the white circle and end at the black circle. Despite its simplicity, the dataset contains a scant number of states wherein multiple actions have been taken. Within this specific framework, the notion of state reachability is straightforward: any state that is just one block away from the current state is considered reachable. Formally the state reachability is, $\mathcal{SR}_{\mathcal{M}}((x,y))=\{(x+1,y),(x-1,y),(x,y+1),(x,y-1)\}$. Although this definition includes the wall states as reachable, SCQL can only perform updates where $s^{\prime}$ is reachable and in the dataset, thus the walls are avoided. In other words, a state is considered reachable from the current state if it is one block away in any of the four cardinal directions (left, right, up, or down). This simple definition of state reachability is sufficient for demonstrating the advantages of SCQL in this illustrative example.
Applying BCQL with training as per Eq. equation 3 and policy extraction as per
$\pi(s)=\max_{a\text{ s.t. }(s,a)\in\mathcal{D}}Q(s,a),$ (8)
we obtain the policy depicted in Figure 2b, where arrows indicate policy direction. Conversely, for SCQL, training via Eq. equation 6 and policy training through Eq. equation 7 yields the policy in Figure 2c. For clarity, if no optimal movement can be found in the state due to no information in the dataset, it is left blank. SCQL is able to leverage the sparse dataset more efficiently, enabling the agent to reach the gold star from any dataset state position. BCQL, however, succeeds in only 11 of the 57 unique dataset states, which is equivalent to the performance of behavioural cloning. SCQL’s advantage stems from its ability to use state adjacency information for trajectory stitching, enabling the agent to construct optimal trajectories even in sparse datasets. By contrast, BCQL’s reliance on previously observed state-action pairs limits its performance when faced with a limited dataset. In many practical settings state reachability cannot be naturally assumed from the environment. As a result, in the following sections, a deep learning implementation of SCQL is defined where state reachability is learned from the dataset.
@@ 4 Practical implementation: StaCQ
In this section, we aim to provide a practical implementation of SCQL. Through this method we seek to benefit from the state-constrained framework, where Q-values are updated on all state and reachable next state pairs; and the policy uses the reachable states to find a more diverse, safe and closer to optimal policy. This requires a practical method to learn state reachability, as well as an actor-critic method that exploits the benefits of the state reachability.
As a result, we introduce StaCQ (State-Constrained deep QSS-learning), our particular deep learning implementation of SCQL. We present a simple, yet effective, approach to learning state reachability, using dynamics models of the environment. StaCQ is a policy constraint method that regularises the policy towards the best reachable next state in the dataset. Our specific implementation learns a QSS-function that is aligned with the underlying theory. However, state-constrained methods can be adapted from the theory in various ways, similar to how many techniques modify the batch constraint in BCQ (Fujimoto et al., 2019). For instance, alternative state-constrained methods might incorporate different policy regularisation techniques, such as those based on divergence measures or constraint violation penalties, or explore different approaches to estimating state reachability using available data and domain knowledge. The purpose here is to illustrate one specific implementation of the broader state-constrained framework, this implementation can be extended or modified in several ways.
@@ 4.1 Estimating state reachability
Central to the state-constrained approach is the concept of state reachability, as per Definition 3.1. Because the environments used in our benchmarks do not provide this information, we need to estimate reachability from the dataset. In our implementation, we consider $\hat{s}^{\prime}$ reachable to $s$ if we can predict the action that reaches $\hat{s}^{\prime}$. For this, we introduce a forward dynamics model, $f_{\omega_{1}}(s,a)$, and an inverse dynamics model, $I_{\omega_{2}}(s,s^{\prime})$. Both models are implemented as neural networks and trained via supervised learning on triplets $(s,a,s^{\prime})\sim\mathcal{D}$. The loss function for the forward model is given by
@@
Consistent with prior research, we adopt an ensemble strategy to account for the epistemic uncertainty (uncertainty in model parameters) (Buckman et al., 2018; Chua et al., 2018; Janner et al., 2019; Argenson & Dulac-Arnold, 2020; Yu et al., 2020; 2021). Each model within the ensemble possesses a distinct set of parameters. The ensemble’s final predictions for $s^{\prime}$ and $a$ are computed by taking the average of their respective outputs. It is important to note that these models are relatively simple, and incorporating more complex architectures, such as those proposed in Zhang et al. (2021), could significantly improve both model predictions and state reachability estimates.
Using these models, we propose an estimator of state reachability, denoted by $\widehat{\mathcal{SR}}_{\mathcal{M}}$, as:
@@
This criterion suggests that $\hat{s}^{\prime}$ is reachable from $s$ only if an action can be predicted that transitions the agent from $s$ to $\hat{s}^{\prime}$. Our practical implementation of state reachability, shown in Eq. equation 11, aims to satisfy Definition 3.1. Without access to the true environment dynamics, we must model these transitions using available data. We use the $L^{\infty}$-norm between observed and predicted states to ensure reachability is determined uniformly across all state dimensions. While other norms, such as $L^{1}$ and $L^{2}$, can be used, our experiments showed no significant difference in the policy produced by StaCQ between the $L^{2}$ and $L^{\infty}$ norms. An ablation study on the effects of norm choice and threshold distance is provided in Appendix F.
It is important to note that $\epsilon$, in Eq. equation 11, is a small positive value used to account for potential model inaccuracies, and we set $\epsilon=0.1$ for all datasets. In Appendix B, we also present a method for reducing the computational complexity of this calculation.
@@ 4.2 StaCQ
StaCQ uses an actor-critic framework, where the critic, represented by the $Q(s,s^{\prime})$ value, is trained by minimising the mean square error (MSE) between predicted and true QSS-values. In line with the theory, the QSS-values are updated across all pairs of states and their reachable next states:
$\mathcal{L}_{\theta}=\mathbb{E}_{\begin{subarray}{c}s\sim\mathcal{D}\\ \hat{s}^{\prime}\in\widehat{\mathcal{SR}}_{\mathcal{M}}(s)\end{subarray}}\Biggl{[}\Biggl{(}r(s,\hat{s}^{\prime})+\gamma Q_{\theta^{\prime}}(\hat{s}^{\prime},f_{\omega_{1}}(\hat{s}^{\prime},\pi_{\phi^{\prime}}(\hat{s}^{\prime})))-Q_{\theta}(s,\hat{s}^{\prime})\Biggr{)}^{2}\Biggr{]}.$ (12)
In this equation, the target actor produces an action, $\pi_{\phi^{\prime}}(\hat{s}^{\prime})$, which is passed into the forward model $f_{\omega_{1}}$ before being evaluated by the QSS-function. The specific reward for the reachable pair $(s,\hat{s}^{\prime})$ may be unseen in the dataset, and if the reward function is unknown, it must be approximated using a neural network. Similar to the forward and inverse dynamics models, this can be achieved through supervised learning by minimising the MSE between the predicted and actual rewards:
@@
Past research suggests that this type of reward model is effective at estimating rewards for previously unseen transitions (Hepburn & Montana, 2024). In Eq. equation 12, $\theta^{\prime}$ represents the parameters of the target Q-value, which are incrementally updated towards $\theta$: $\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$, where $\tau$ is the soft update coefficient.
To train the deterministic actor, $\pi:\mathcal{S}\rightarrow\mathcal{A}$, we aim to maximise the current QSS-values while staying close to the best next state in the dataset. Therefore, the loss we minimise is:
$\mathcal{L}_{\phi}=\mathbb{E}_{s\sim\mathcal{D}}\Bigl{[}-\lambda Q_{\theta}(s,f_{\omega_{1}}(s,\pi_{\phi}(s)))+(f_{\omega_{1}}(s,\pi_{\phi}(s)))-\hat{s}^{\prime})^{2}\Bigr{]},$ (14)
where the hyperparameter $\lambda$ controls the level of regularisation. The $\hat{s}^{\prime}$ represents the best reachable next state from $s$ according to the QSS-value,
$\hat{s}^{\prime}=\operatorname*{arg\,max}_{s^{\prime}\in\widehat{\mathcal{SR}}_{\mathcal{M}}(s)}Q_{\theta}(s,s^{\prime}).$ (15)
This is, therefore, a similar policy extraction method to TD3+BC (Fujimoto & Gu, 2021), as both methods use a MSE regulariser. However, StaCQ leverages the state reachability metric, allowing the maximisation to be performed over multiple states, whereas TD3+BC is limited to a single state-action pair.
@@
3:  Pre-train $f_{\omega_{1}}$ $I_{\omega_{2}}$: Eqs.equation 9 & equation 10
@@
6:     Optimise reward function: Eq. equation 13
7:     Optimise QSS-value: Eq. equation 12
8:     Optimise policy: Eq. equation 14
9:     Update target networks: $\theta^{\prime}\leftarrow\tau\theta+(1-\tau)\theta^{\prime}$,$\phi^{\prime}\leftarrow\tau\phi+(1-\tau)\phi^{\prime}$
@@
Similar to batch-constrained offline RL methods, StaCQ aims to balance staying close to the dataset while maximising the Q-value. The hyperparameter $\lambda$ controls the trade-off between producing actions that lead to high-value OOD states and staying close to states in the dataset, specifically the highest value reachable state. Since the forward model is fixed, a small amount of Gaussian noise is added to the policy action, to ensure the policy does not overfit to the forward model, resulting in a more robust policy. The full algorithm is presented in Algorithm 1. Notably, all three constituent models are trained using supervised learning, making them straightforward sub-processes within the overall architecture. The code can be found at https://github.com/CharlesHepburn1/State-Constrained-Offline-Reinforcement-Learning.
5 Related work
Model-free offline RL.
BCQ has emerged as one of the pioneering offline deep RL methodologies (Fujimoto et al., 2019). It posits that having all $(s,a)$ pairs in the MDP enables the construction of an optimal QSA-value by focusing updates only on the $(s,a)$ pairs present in the dataset. As a result, BCQ restricts the policy to these observed $(s,a)$ pairs while allowing minimal perturbations to facilitate marginal improvements. Following BCQ, numerous algorithms have been introduced to develop alternative ways to constrain the policy to seen $(s,a)$ pairs, either by restricting it to the support of the $(s,a)$-distribution (Kumar et al., 2019; Wu et al., 2019; Siegel et al., 2020; Kostrikov et al., 2021a; Zhou et al., 2021; Brandfonbrener et al., 2021) or directly limiting it to the $(s,a)$ pairs (Fujimoto & Gu, 2021).
In contrast to policy constraints, some methods tackle distributional shifts by pessimistically evaluating the value function on OOD $(s,a)$ pairs (Kumar et al., 2020; Yu et al., 2021; An et al., 2021). There is also a subset of offline RL methods that combines both pessimistic value evaluation and policy constraints (Dadashi et al., 2021; Beeson & Montana, 2024). Alternatively, implicit Q-learning (IQL) (Kostrikov et al., 2021b) and policy-guided offline RL (POR) (Xu et al., 2022) aim to learn an optimal value function through expectile regression, with POR estimating a state-value function and IQL estimating an action-value function. Both methods learn the optimal policy (referred to as the guide policy in POR) via advantage-weighted regression. All of these approaches adopt the batch-constrained method, prioritising explicit $(s,a)$ pairs or $(s,s^{\prime})$ pairs found in the dataset over unobserved actions. Our state-constrained method, however, relaxes the batch-constrained objective by requiring constraints only on reachable states in the dataset rather than explicit $(s,a)$ pairs.
Model-based offline RL.
Model-based methods learn the dynamics of the environment to support policy learning (Sutton, 1991; Janner et al., 2019). There are generally two main approaches to model-based offline RL. The first approach learns a pessimistic model of the environment and performs rollouts with this augmented model, effectively increasing the dataset size (Yu et al., 2020; Kidambi et al., 2020; Rigter et al., 2022). The second approach uses the models for planning, allowing the agent to look ahead during evaluation to determine the optimal path (Argenson & Dulac-Arnold, 2020; Zhan et al., 2021; Janner et al., 2022; Diehl et al., 2021). In contrast, StaCQ uses models to implement our concept of state reachability; thus, StaCQ does not generate new states or involve planning during evaluation.
State reachability.
The state-constrained approach relies heavily on the concept of state reachability. Prior studies on state reachability (Hepburn & Montana, 2022; 2024) have employed a Gaussian distribution to determine the likelihood of stitching from $s$ to $s^{\prime}$. In contrast, our method treats a state as either reachable or not, avoiding the use of continuous probability distributions. State reachability is closely related to state similarity, a well-explored concept in the literature (Zhang et al., 2020; Agarwal et al., 2021; Le Lan et al., 2021). One way to measure state similarity is through bisimulation metrics, which are based on the environment’s dynamics (Ferns et al., 2012). Traditional bisimulation methods often require full state enumeration (Chen et al., 2012; Bacci et al., 2013a; b), leading to the development of more scalable pseudometric approaches (Castro, 2020). While incorporating a pseudometric into our state reachability framework could be a potential avenue, we leave this for future work. Our current state reachability metric is simple, intuitive, and grounded in the core definition.
Trajectory stitching.
A key feature of offline RL is its ability to stitch together trajectories (Kostrikov et al., 2021b), i.e., combining previously observed trajectories to create a novel one that completes the task. Imitation learning methods, such as behavioural cloning (Pomerleau, 1988; 1991), struggle with stitching due to their inability to distinguish between optimal and sub-optimal states. Similarly, the Decision Transformer (DT) (Chen et al., 2021) formulates the offline RL problem as a supervised learning task using a goal-conditioned policy. While DT exhibits weak stitching capabilities, efforts have been made to address this by integrating offline RL principles (Yamagata et al., 2023; Wu et al., 2023). Combining DT with StaCQ’s policy extraction approach could yield further improvements in stitching. However, our current work focuses on a straightforward implementation of state-constrained offline RL.
@@ 6 Experimental results
We evaluate StaCQ against several model-free and model-based baselines on the D4RL benchmarking datasets from the OpenAI Mujoco tasks (Todorov et al., 2012; Fu et al., 2020). The model-free baselines we compare against are BCQ (Fujimoto et al., 2019), TD3+BC (Fujimoto & Gu, 2021), and IQL (Kostrikov et al., 2021b). BCQ is the foundational batch-constrained offline RL method on which many subsequent methods are based, making it the most direct theoretical comparison. TD3+BC is the most similar to StaCQ in terms of implementation, as both use a BC-style regularisation process. IQL, a current SOTA model-free method, provides a strong comparison point.
The model-based baselines include MBTS (Hepburn & Montana, 2024), Diffuser (Janner et al., 2022), and RAMBO (Rigter et al., 2022). MBTS employs a data augmentation strategy with a different state reachability approach compared to StaCQ. Diffuser is a planning-based method that uses a diffusion model, making its dynamics model more complex than StaCQ’s. RAMBO, a SOTA model-based method, performs rollouts using a pessimistic dynamics model of the environment. Although StaCQ does not use models for planning or rollouts, we include these model-based algorithms for comparison.
Additionally, we devised a one-step version of StaCQ (Brandfonbrener et al., 2021), detailed in Appendix D, to showcase the flexibility of the state-constrained framework.
@@
Hopper Rand $6.2$ $7.6$ $8.5$ $\mathbf{21.6}$ $17.5\pm 12.8$ $7.5\pm 0.4$
Med-Rep $22.5$ $51.0$ $60.9$ $\mathbf{94.7}$ $50.2$ $93.6$ $\mathbf{96.6}$ $\mathbf{99.1\pm 1.3}$ $\mathbf{99.4\pm 0.6}$
Med $56.8$ $60.9$ $59.3$ $66.3$ $64.3$ $74.3$ $92.8$ $\mathbf{100.2\pm 3.0}$ $93.3\pm 3.2$
Med-Exp $54.2$ $85.9$ $98.0$ $91.5$ $94.8$ $103.3$ $83.3$ $\mathbf{111.9\pm 0.2}$ $92.1\pm 9.2$
@@
Med $39.4$ $73.7$ $83.7$ $78.3$ $78.8$ $79.6$ $86.9$ $\mathbf{92.2\pm 3.0}$ $85.7\pm 10.4$
Med-Exp $90.5$ $94.5$ $\mathbf{110.1}$ $\mathbf{109.6}$ $108.8$ $106.9$ $68.3$ $\mathbf{116.2\pm 1.5}$ $107.6\pm 8.5$